OpenStudy (anonymous):

can someone help me to derivate the following function? f(x)=ln(3x)

OpenStudy (anonymous):

ln' ( u) = u'/ u u = 3x --> u' = ...

OpenStudy (zarkon):

\[\ln(3x)=\ln(3)+\ln(x)\] then differentiate